OpenStudy (anonymous):

Jenny and Kalim collect rare coins. Jenny has 11 fewer coins than Kalim, who has "y" coins. Jenny added 5 coins to her collection. Which of the following is an expression for the number of coins that Jenny has? 6 y+6 y+15 2y-15 y-6

OpenStudy (anonymous):

Y-6. If Jenny had 11 coins fewer than Kalim (Y) then the equation would be: J=Y-11. Jenny then added 5 coins: J=Y-11+5 Simplify to get: J=Y-6.